# Data Generation for Stability Studies of Power Systems with High Penetration of Inverter-Based Resources

### v1

@@ III-B Analysis of The Generated Datasets
To validate the performance of the proposed tool, datasets are generated under different parameter settings. The resulting data are then analyzed in terms of their composition and how they evolve during the execution of the data generation process. Finally, an Extreme Gradient Boosting (XGB) model [xgboost_python] is trained, and its accuracy is evaluated.
@@
Two datasets are created, and their main difference lies in whether sensitivity information is used during the data generation process. In Test #1, sensitivity is not used; in this case, $P_{SG}$ and $P_{IBR}$ are always split, following the approach in [rossi2022data]. In Test #2, sensitivity is used to determine along which dimension the split should be performed, selecting the one that most strongly affects the stability outcome before splitting each cell. This aims to demonstrate the necessity of incorporating sensitivity information for larger systems and the scalability improvements enabled by the proposed tool. The other difference in the execution of the data generation process concerns the maximum exploration depth allowed. All parameter settings are summarized in Table II. The main characteristics of the resulting datasets (namely, the percentage of feasible, discarded, and stable instances) are also reported.
TABLE II: Summary of parameter settings and datasets composition
@@ Appendix A Dynamic Equivalents of IBRs Aggregation
This section presents the admittance-based frequency scan analysis carried out to verify the dynamic equivalence of the IBRs aggregation employed in the data generation tool. The comparison between the admittance-based frequency scans of the base scenarios (with individually modeled IBRs) and the corresponding aggregated scenarios is performed under consistent operating conditions. The aggregated converter is assigned a nominal power equal to the sum of the nominal powers of the individual converters it represents, and its operating point corresponds to the total active power output of those converters. Additionally, the voltage at the bus to which the IBRs are connected, as well as the voltage at the bus representing the Thevenin equivalent of the remainder of the grid, are kept equal across both scenarios. Figure 12 illustrates two base scenarios along with their corresponding aggregated equivalents. Figures 13(a) and 13(b) present the comparison between the respective admittance-based frequency scans, demonstrating a perfect match between the responses obtained in the base scenarios and those derived from the equivalent aggregated representations.
@@
Figure 13: Admittance frequency-scans comparison.
@@ Appendix B Tools for Power System Operation and Stability Analysis
To assess the small-signal stability of the sampled operating points and associated  IBRs control settings, conventional power system stability analysis tools are employed. These tools typically rely on the linearized state-space representation of the system. The eigenvalue analysis of the system’s state matrix provides insight into system stability. The linearization process is performed around an equilibrium point, which can be obtained either through time-domain simulations or PF calculations. In this work, the latter approach is preferred due to its computational efficiency. It is important to note that, despite the sampling strategy being designed to respect key relationships, it is still likely that many sampled operating points do not satisfy the system’s feasibility constraints. This is primarily due to the voltage profiles that, even if sampled considering the existing relation between adjacent buses, may lead to unrealistic or infeasible steady-state conditions, such as reactive power injections from generators that violate their capability curves. Additional sources of infeasibility include violations of line thermal ratings, voltage magnitude limits, and generators power factor. To ensure that only feasible operating points are analyzed, these constraints must be accounted for. Therefore, instead of performing a standard PF calculation, an OPF problem is solved. The objective function minimizes the squared difference between the sampled active power set point of each generator and its resulting active power injection. This formulation allows the system to adjust controllable variables within operational limits to satisfy all relevant constraints, ensuring that the resulting operating point is both physically realistic and technically feasible. To implement these calculations in a Python environment, the feasible operating point is computed using the VeraGrid [veragrid] library. The AC-OPF implemented in VeraGrid, which considers voltage, current, and generator power factor constraints as described in [alegre2024integration], has been adapted to incorporate the objective function described above. For the small-signal stability analysis, the Python version of the STAMP tool [arevalo2025matlab] is used, which is available at [stampy].
Acknowledgment

### v2

@@ 3.2 Analysis of The Generated Datasets
To validate the performance of the proposed tool, datasets are generated under different parameter settings. The resulting data are then analyzed in terms of their composition and how they evolve during the execution of the data generation process. Finally, an Extreme Gradient Boosting (XGB) model [22] is trained, and its accuracy is evaluated.
@@
Two datasets are created, and their main difference lies in whether sensitivity information is used during the data generation process. In Test #1, sensitivity is not used; in this case, $P_{SG}$ and $P_{IBR}$ are always split, following the approach in [16]. In Test #2, sensitivity is used to determine along which dimension the split should be performed, selecting the one that most strongly affects the stability outcome before splitting each cell. This aims to demonstrate the necessity of incorporating sensitivity information for larger systems and the scalability improvements enabled by the proposed tool. The other difference in the execution of the data generation process concerns the maximum exploration depth allowed. All parameter settings are summarized in Table 2. The main characteristics of the resulting datasets (namely, the percentage of feasible, discarded, and stable instances) are also reported.
Table 2: Summary of parameter settings and datasets composition
@@ Appendix A Dynamic Equivalents of IBRs Aggregation
This section presents the admittance-based frequency scan analysis carried out to verify the dynamic equivalence of the IBRs aggregation employed in the data generation tool. The comparison between the admittance-based frequency scans of the base scenarios (with individually modeled IBRs) and the corresponding aggregated scenarios is performed under consistent operating conditions. The aggregated converter is assigned a nominal power equal to the sum of the nominal powers of the individual converters it represents, and its operating point corresponds to the total active power output of those converters. Additionally, the voltage at the bus to which the IBRs are connected, as well as the voltage at the bus representing the Thevenin equivalent of the remainder of the grid, are kept equal across both scenarios. Figure 13 illustrates two base scenarios along with their corresponding aggregated equivalents. Figures 14(a) and 14(b) present the comparison between the respective admittance-based frequency scans, demonstrating a perfect match between the responses obtained in the base scenarios and those derived from the equivalent aggregated representations.
@@
Figure 14: Admittance frequency-scan comparison.
@@ Appendix B Tools for Power System Operation and Stability Analysis
To assess the small-signal stability of the sampled operating points and associated  IBRs control settings, conventional power system stability analysis tools are employed. These tools typically rely on the linearized state-space representation of the system. The eigenvalue analysis of the system’s state matrix provides insight into system stability. The linearization process is performed around an equilibrium point, which can be obtained either through time-domain simulations or PF calculations. In this work, the latter approach is preferred due to its computational efficiency. It is important to note that, despite the sampling strategy being designed to respect key relationships, it is still likely that many sampled operating points do not satisfy the system’s feasibility constraints. This is primarily due to the voltage profiles that, even if sampled considering the existing relation between adjacent buses, may lead to unrealistic or infeasible steady-state conditions, such as reactive power injections from generators that violate their capability curves. Additional sources of infeasibility include violations of line thermal ratings, voltage magnitude limits, and generators power factor. To ensure that only feasible operating points are analyzed, these constraints must be accounted for. Therefore, instead of performing a standard PF calculation, an OPF problem is solved. The objective function minimizes the squared difference between the sampled active power set point of each generator and its resulting active power injection. This formulation allows the system to adjust controllable variables within operational limits to satisfy all relevant constraints, ensuring that the resulting operating point is both physically realistic and technically feasible. To implement these calculations in a Python environment, the feasible operating point is computed using the VeraGrid [23] library. The AC-OPF implemented in VeraGrid, which considers voltage, current, and generator power factor constraints as described in [24], has been adapted to incorporate the objective function described above. For the small-signal stability analysis, the Python version of the STAMP tool [25] is used, which is available at [26].
Declaration of Generative AI and AI-assisted technologies in the writing process
During the preparation of this work, the authors used the tool ChatGPT developed by OpenAI to improve the readability. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the publication’s content.
Acknowledgments